# Steering LLM Reasoning Through Bias-Only Adaptation

### v2

@@ Abstract
We show that training a single $d$-dimensional steering vector per layer with reinforcement learning, while freezing all base weights, matches the accuracy of fully RL-tuned reasoning models on mathematical-reasoning tasks. On an 8 billion-parameter model this adds only $\approx 0.0016\%$ additional parameters and reproduces performance across a range of base models and mathematical-reasoning benchmarks. These results tighten the upper bound on the parameter budget required for high-level chain-of-thought reasoning, indicating that millions of adapter weights are unnecessary. The minimal trainable footprint reduces optimizer memory and inter-GPU communication, lowering the overall cost of fine-tuning. Moreover, a logit-lens analysis shows that the learned vectors amplify coherent token directions, providing clearer insight into the model’s internal computations.
@@
Viacheslav Sinii1, Alexey Gorbatovski1,2, Artem Cherepanov1,3, Boris Shaposhnikov1, Nikita Balagansky1,2, Daniil Gavrilov1, 1T-Tech, 2Moscow Institute of Physics and Technology, 3Central University, Correspondence: v.siniy@tbank.ru
@@ 2 Related Work
Our work is partly motivated by the recent results on reasoning training only amplifying the reasoning behaviors already present in the base models (Wang et al., 2025; Ye et al., 2025; Shao et al., 2025; Liu et al., 2025). If such a minimal intervention is able to recover reasoning performance, this adds as an evidence to such a claim. Our aim is to take the study further by identifying what specific parts of the model are holding these behaviours.
@@ 3.3 Training and Evaluation Setup
We report results on six math benchmarks: AIME24/25, AMC23, MATH500 (Hendrycks et al., 2021a), MinervaMath (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). For MATH500, MinervaMath, and OlympiadBench we report Pass@1; for AIME24/25 and AMC23 we report Avg@32 due to their smaller size. Base models decode greedily, whereas trained models use sampling with $\tau=1.0$ following Zeng et al. (2025). Evaluation context length is 4K and 32K for Qwen2.5-Math-7B and other models respectively. All metrics are averaged over three evaluation seeds. Unless noted otherwise (e.g., Table˜1), figures and tables show the mean score across the six benchmarks. When available, we include numbers from SimpleRL-Zoo (Zeng et al., 2025), Oat-Zero (Venhoff et al., 2025), Open-Reasoner (Hu et al., 2025), and R1-Distill (Guo et al., 2025) for context. Implementations use transformers (Wolf et al., 2019), the vllm inference engine Kwon et al. (2023), and Math-Verify [^1] https://github.com/huggingface/Math-Verify for reward assignment.
@@ 5 Conclusion
In this paper, we have demonstrated that training lightweight steering vectors alone can recover the reasoning performance of fully-tuned models on standard mathematical benchmarks. This result carries two important implications. First, steering vector training offers a highly parameter-efficient and cost-effective alternative: only a small set of layer-wise bias terms must be learned, drastically reducing storage and communication time requirements. Second, it isolates a small, interpretable set of parameters that capture the effects of reasoning training, simplifying the study of the mechanisms of models’ reasoning abilities.
@@ Appendix G Training Cost Savings
Table 5: Resource-efficiency comparison of full fine-tuning versus steering across three model sizes.

### v3

@@ Abstract
We show that training a single $d$-dimensional steering vector per layer with reinforcement learning, while freezing all base weights, matches the accuracy of fully RL-tuned reasoning models on mathematical-reasoning tasks. On an 8 billion-parameter model this adds only $\approx 0.0016\%$ additional parameters and reproduces performance across a range of base models and mathematical-reasoning benchmarks. These results tighten the upper bound on the parameter budget required for high-level chain-of-thought reasoning, indicating that millions of adapter weights are unnecessary. The minimal trainable footprint reduces optimizer memory and inter-GPU communication, lowering the overall cost of fine-tuning. Moreover, a logit-lens analysis shows that the learned vectors amplify coherent token directions, providing clearer insight into the model’s internal computations. The code is available at https://github.com/corl-team/steering-reasoning.
@@
Viacheslav Sinii1, Alexey Gorbatovski1, Artem Cherepanov1,2, Boris Shaposhnikov1, Nikita Balagansky1, Daniil Gavrilov1, 1T-Tech, 2Central University, Correspondence: v.siniy@t-tech.dev
@@ 2 Related Work
Our work is partly motivated by the recent results on reasoning training only amplifying the reasoning behaviors already present in the base models (Wang et al., 2025; Ye et al., 2025; Shao et al., 2025; Liu et al., 2025a). If such a minimal intervention is able to recover reasoning performance, this adds as an evidence to such a claim. Our aim is to take the study further by identifying what specific parts of the model are holding these behaviours.
@@ 3.3 Training and Evaluation Setup
We report results on six math benchmarks: AIME24/25, AMC23, MATH500 (Hendrycks et al., 2021a), MinervaMath (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). For MATH500, MinervaMath, and OlympiadBench we report Pass@1; for AIME24/25 and AMC23 we report Avg@32 due to their smaller size. Base models decode greedily, whereas trained models use sampling with $\tau=1.0$ following Zeng et al. (2025). Evaluation context length is 4K and 32K for Qwen2.5-Math-7B and other models respectively. All metrics are averaged over three evaluation seeds. Unless noted otherwise (e.g., Table˜1), figures and tables show the mean score across the six benchmarks. When available, we include numbers from SimpleRL-Zoo (Zeng et al., 2025), Oat-Zero (Liu et al., 2025b), Open-Reasoner (Hu et al., 2025), and R1-Distill (Guo et al., 2025) for context. Implementations use transformers (Wolf et al., 2019), the vllm inference engine Kwon et al. (2023), and Math-Verify [^1] https://github.com/huggingface/Math-Verify for reward assignment.
@@ 5 Conclusion
In this paper, we have demonstrated that training lightweight steering vectors alone can recover the reasoning performance of fully‐tuned models on standard mathematical benchmarks. This result carries two important implications. First, steering vector training offers a highly parameter‐efficient and cost‐effective alternative: only a small set of layer‐wise bias terms must be learned, drastically reducing storage and communication time requirements. Second, it isolates a small, interpretable set of parameters that capture the effects of reasoning training, simplifying the study of the mechanisms of models’ reasoning abilities.
@@ Appendix G Training Cost Savings
Table 5: Resource‐efficiency comparison of full fine-tuning versus steering across three model sizes.